OpenStudy (anonymous):

WILL FAN + MEDAL Luella took out an unsubsidized student loan of $37,000 at a 4.8% APR, compounded monthly, to pay for her last six semesters of college. If she will begin paying off the loan in 33 months with monthly payments lasting for 20 years, what will be the amount of her monthly payment? A.$271.56 B.$240.11 C.$238.04 D.$273.92

OpenStudy (anonymous):

i really need help here....

OpenStudy (amistre64):

compound the 37k for 33 months, what do we get for balance to pay off?

OpenStudy (anonymous):

oh and amistre? ur formula didn't really help me..... :P

OpenStudy (anonymous):

First find out how much she owes (i.e. how much interest has accrued in addition to the initial loan amount) using the compound interest formula: \[A=P\left(1+\frac{r}{n}\right)^{nt}\] (if I'm not mistaken) Then I believe the present value formula should be used to figure out payment size.

OpenStudy (anonymous):

oo

OpenStudy (amistre64):

i worked it out on the wolf and linked it ...

OpenStudy (anonymous):

alright fine lol

OpenStudy (anonymous):

u get an award for trying anyway xD

OpenStudy (amistre64):

it even works for this one :)

OpenStudy (amistre64):

\[P=37000k^{33}~k^{12*20}~\frac{1-k}{1-k^{12*20}},~k=1+.048/12\] http://www.wolframalpha.com/input/?i=37000k%5E%2833%29k%5E%2812*20%29+%281-k%29%2F%281-k%5E%2812*20%29%29%2Ck%3D1%2B.048%2F12

OpenStudy (anonymous):

how should i input that into my calculator?

OpenStudy (amistre64):

depends on the calculator you have; i use a ti83 .... ive never used a financial calculator tho

OpenStudy (anonymous):

i use a ti84 silver edition

OpenStudy (amistre64):

you know how to store things into the alpha characters?

OpenStudy (anonymous):

on tvm?

OpenStudy (anonymous):

or the fracions?

OpenStudy (amistre64):

i like using K as a clean up for the compounding mess: 1+.048/12, STO \(\to\), alpha, K ; the "(" key is "K" 37000, alpha, K, ^, (33), STO \(\to\), alpha, B then im ready to run the setup

OpenStudy (amistre64):

B K^(12*20) (1-k) \(\div\) (1-K^(12*20))

OpenStudy (anonymous):

lol its a little too confusing here

OpenStudy (amistre64):

then id just stick with the wolfs output :)

OpenStudy (anonymous):

well how do you get the answers?! its vital!!!!

OpenStudy (anonymous):

oh wait.... >_<

OpenStudy (anonymous):

it was posted on htere :P

OpenStudy (anonymous):

hallelu-friggin-jah finally... on to question 6

OpenStudy (anonymous):

5*

OpenStudy (amistre64):

id have to go thru some more confusing things .... like this: lets start with a begining loan amount Bo\[B_o=B_o\]and take this thru each month with a compounding variable ... k \[B_1=B_ok\] \[B_2=B_o~k~k=B_ok^2\] \[B_3=B_ok^3\] do this for 33 months to get:\[B_{33}=B_ok^{33}\]then we can use that amount to start paying off \[B_{34}=B_ok^{34}-P\] \[B_{35}=B_ok^{35}-Pk-P\] \[B_{36}=B_ok^{36}-Pk^2-Pk-P\] this goes on and on and forms a geometric sum of the payments in terms of k \[B_{n}=B_ok^{d+n}-P\frac{1-k^n}{1-k}\] when Bn = 0 the loan is paid off \[0=B_ok^{d+n}-P\frac{1-k^n}{1-k}\] so solving for P is the payment amount \[B_ok^{d+n}=P\frac{1-k^n}{1-k}\] \[B_ok^{d+n}\frac{1-k}{1-k^n}=P\]

OpenStudy (amistre64):

soo, with the specifics given \[P=37000\left(1+\frac{.048}{12}\right)^{33+12(20)}~\frac{1-(1+\frac{.048}{12})}{1-(1+\frac{.048}{12})^{12(20)}}\]

OpenStudy (anonymous):

...